# GeoGuess: Multimodal Reasoning based on Hierarchy of Visual Information in Street View

### v1

@@ 4.3. Visual Clue Detection
(4) $=\cup_{j=1}^{im}\text{OOD}(I_{j},prompt)$
Based on analysis of high-frequency nouns in the explanation $E$ and knowledge snippets $S$, we design a comprehensive prompt to capture all potentially identifiable object. We use Grounding DINO(Liu et al., 2024b) as our open-vocabulary object detector in this stage.
4.4. Multimodal Knowledge Retrieval
Since GeoGuess is an expert-level task even for humans, relying on the inherent commonsense knowledge in MLLMs obtained from large-scale pretraining, or on reasoning techniques like (Wei et al., 2022; Yao et al., 2023), is insufficient for generating accurate answers. To alleviate potential hallucinations produced by LLM, we have to explicitly incorporate external geographical knowledge into the reasoning process. Accordingly, in the second stage, we aim to retrieve informative knowledge snippets from GeoKnowledge in a multimodal way to assist the reasoning process, inspired by (Lewis et al., 2020).
@@ 4.4. Multimodal Knowledge Retrieval
(5) $\mathbf{S^{*}}$ =$=\{S_{i_{1}^{*}},S_{i_{2}^{*}},S_{i_{3}^{*}}\}$
(6) $\mathbf{V^{*}}$ =$=\{V_{j_{1}^{*}},V_{j_{2}^{*}},V_{j_{3}^{*}}\}$
(7) $i_{1}^{*},i_{2}^{*},i_{3}^{*},j_{1}^{*},j_{2}^{*},j_{3}^{*}$ $=\arg\operatorname{Top\text{-}3}_{i,j}\frac{\text{VE}(KI_{i})\cdot\text{VE}(V_{j})}{|\text{VE}(KI_{i})||\text{VE}(V_{j})|}$
In Equation 7, $KI$ denotes an image of country-specific object from GeoKnowledge, and $V$ represents detected visual clues obtained in the first stage. To retrieve the top 3 similar image in GeoKnowledge, we employ cosine similarity. A visual encoder, denoted as VE, is used to embed images into vector representations. In our work, we adopt CLIP-ViT (Radford et al., 2021) as the visual encoder with an embedding dimension of 768.
4.5. Reasoning Generation
@@
After constructing both image and text prompt, we feed them into a pre-trained MLLM to perform reasoning and generate final answer. Due to the varying solutions of images used in the image prompt, we choose Qwen2VL (Wang et al., 2024) as our base model which is good at understanding flexible visual inputs.
4.6. Training Policy
@@
(8) $\text{Loss}=-\sum_{t=1}^{T}P(x_{t}|x_{1},\dots,x_{t-1})$
Since our dataset is relatively small compared with pre-training corpus, we incorporate LoRA (Hu et al., 2022) to fine-tune only a small subset of parameters. This helps reduce the demand for training resources, the risk of overfitting, and enables faster convergence.
5. Experiments
@@
5.1. Experiments Setup
SightSense utilizes GroundingDINO (Liu et al., 2024b) as the open-vocabulary object detector in the visual clue detection stage, CLIP-ViT (Radford et al., 2021) as the visual encoder in the multimodal knowledge retrieval stage and Qwen2VL-7B-Instruct (Wang et al., 2024) as the base reasoning model in reasoning generation stage. We finetune only a subset of parameters in Qwen2VL using LoRA (Hu et al., 2022), while keeping the parameters of models in GroundingDINO and CLIP-ViT frozen. For finetuning details, we adopt Adam optimizer (Kingma and Ba, 2014) to optimize the training objective and use LinearLR to adjust the learning rate. All training work was completed on NVIDIA H100 PCIe.
5.2. Evaluation Metrics
Each sample in GeoExplain includes a ground-truth geo-coordinates together with several comprehensive reference explanations describing how the correct location can be inferred. We refer to the location and explanations generated by SightSense and other baselines as candidate location and explanations. Evaluation of the GeoGuess task comprises two aspects: (1) the accuracy of the candidate location, and (2) the textual similarity between model-generated candidate explanations and human-written references which is measured by both automatic metrics and human evaluation.
5.2.1. Accuracy of Location
We convert the ground-truth geo-coordinate into structured ground-truth location in the format (COUNTRY, STATE, CITY, STREET) which matches the format of candidate location. In case that some countries may have different names, we apply fuzzy match. Accuracy is computed at each of the four hierarchical levels by measuring the proportion of correct matches over total number of samples. The accuracy can be formulated as Equation 9, where Match is the fuzzy match function, $c_{i}$ and $gt_{i}$ denotes the candidate location and ground-truth location respectively.
(9) $\text{Acc}=\frac{\sum_{i=1}^{n}\text{Match}(c_{i},gt_{i})}{n}$
5.2.2. Explanation: Automatic Metrics
In order to measure the similarity between candidate and reference explanations and the comprehensiveness of explanation, we employ two aspects of automatic metrics.
  • Relevance: We use (1) BLEU (Papineni et al., 2002), (2) Rogue-L (Lin, 2004), (3) CIDEr (Vedantam et al., 2015), (4) Meteor (Banerjee and Lavie, 2005) and (5) BERTScore (Zhang et al., 2019) to evaluate the relevance between candidate and reference explanations.
  • Diversity: We use (1) Dist-1 and (2) Dist-2 (Li et al., 2016) to measure the diversity of candidate explanations themselves.
@@ 5.2.2. Explanation: Automatic Metrics
P+VC Qwen2VL-7B-VC 3.00 $\uparrow$ 0.90 $\downarrow$ 18.85 $\uparrow$ 1.03 $\downarrow$ 28.29 $\downarrow$ 84.98 $\uparrow$ 5.02 $\downarrow$ 17.70 $\downarrow$
@@
P+VC+K Qwen2VL-7B-KVC 5.44 $\uparrow$ 2.03 $\uparrow$ 20.00 $\uparrow$ 1.12 $\uparrow$ 29.35 $\uparrow$ 85.39 $\uparrow$ 4.58 $\downarrow$ 15.37 $\downarrow$
GPT-4o-mini-KVC 2.45 $\uparrow$ 0.40 $\downarrow$ 13.32 $\uparrow$ 0.37 $\uparrow$ 22.42 $\downarrow$ 82.88 -- 7.49 $\uparrow$ 28.16 $\downarrow$
@@
Table 2. Automatic evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines and underlying number denotes the highest accracy for each type. $\uparrow$, $\downarrow$ and -- indicates how the accuracy change for corresponding models when introducing more modules.
5.2.3. Explanation: Human Evaluation
@@
5.3. Baselines
We compare SightSense with several state-of-the-art baselines in multimodal reasoning. We validate the effectiveness of hierarchical visual information in the reasoning process even for other baselines by transplanting the first two stages of our pipeline into them. This enables a fair comparison and highlights the generalizability of our approach. Furthermore, we classify the baselines into three different types according to information they receive during generating explanations.
•  Panorama Only: only takes the thumbnail image of the street view panorama as reasoning inputs. In this type, we use SightSense_wKVC which is a simplified version of SightSense, removing the Visual Clue Detection and Knowledge Retrieval stage. For baseline models, we selected well-established and widely recognized state-of-the-art models with comparable model sizes.: (1) vanilla Qwen2VL-7B without finetuning; (2) GPT-4o-mini and (3) DeepSeek-VL2-tiny. These models serve as strong and representative baselines for evaluating the performance of our approach.
•  Panorama + Visual Clue: reasons on hierarchical visual information by integrating the panorama and detected visual clues into reasoning. Similarly, we finetune SightSense_wK which removes only the second stage from SightSense. Meanwhile, we transplant the Visual CLue Detection stage into other baselines. We denote them as (1) Qwen2VL-7B-VC; (2)GPT-4o-mini-VC and (3) DeepSeek-VL2-tiny-VC.
•  Panorama + Visual Clue + Knowledge: reasons on hierarchical visual information and external knowledge. We use SightSense with all components enabled. Accordingly, we refer to the baseline models with all three modules as (1) Qwen2VL-7B-KVC; (2)GPT-4o-mini-KVC and (3) DeepSeek-VL2-tiny-KVC.
5.4. Main Results
@@
P+VC Qwen2VL-7B-VC 16.62 $\downarrow$ 1.38 $\uparrow$ 0.00 -- 0.00 --
GPT-4o-mini-VC 11.38 $\uparrow$ 0.92 $\uparrow$ 0.00 -- 0.00 --
DeepSeek-VL2-tiny-VC 29.79 $\uparrow$ 5.32 $\uparrow$ 0.09 -- 0.00 --
@@
P+VC+K Qwen2VL-7B-KVC 25.80 $\uparrow$ 4.91 $\uparrow$ 0.10 $\uparrow$ 0.09 $\uparrow$
@@
Table 3. Accuracy of candidate location in %. Denotations are the same as in Table 2
@@
Table 4. Human Evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines.
@@
Figure 4. A case study of explanation generated by SightSense, Qwen2VL-7B, GPT-4o-mini and DeepSeek-VL2-tiny. Given image is a thumbnail of panoramas set and the groundtruth is on the right hand side.
@@
•  Accuracy of Candidate Location: From Table 3, SightSense substantially outperforms all other baselines across all four levels, achieving notable gains of 27% and 18% on the Country and State levels respectively. Meanwhile, within each type, the simplified version of SightSense also almost achieves the best performance. This verifies the effectiveness of the architecture and training objective of SightSense and highlights its advantage in addressing the GeoGuess task. From the comparison across the same model in different types, e.g., Qwen2VL-7B, Qwen2VL-7B-VC and Qwen2VL-7B-KVC, introducing Visual Clue Detection module results in 8 accuracy increases, 2 decreases and 6 ties. Similarly, introducing Multimodal Knowledge Retrieval module leads to 15 increases and 1 decrease. The comparison demonstrates that integrating hierarchical visual information and external knowledge consistently improves location prediction accuracy. The consistent accuracy gains across all four models further verify the robustness and generalizability of our method.
@@
•  Automatic Metric of Candidate Explanation: We observe SightSense series still achieves the best performance regarding explanation similarity evaluation across all types in 2. Explanations generated by SightSense show the highest similarity with the reference explanations. Furthermore, among all baseline models, the origin version of SightSense achieves the highest relevance score. The diversity metric is primarily dominated by DeepSeek-VL2 which means SightSense tends to generate a fixed template learned from the GeoExplain dataset. As with location accuracy analysis, we also notice significant increase in similarity metric when introducing Visual Clue Detection and Multimodal Knowledge Retrieval module, proving the effectiveness of our approach.
@@
•  Human Evaluation of Candidate Explanation: Table 4 demonstrates the result of human evaluation. Due to large workload, we ask the volunteers to rate only SightSense and three base models. Except for Fluency, SightSense achieves the highest score across all other metrics. This indicates that the explanation generated by SightSense provide the strongest evidence, the broadest and the most relevant geographical knowledge, and the most complete reasoning chain. These improvements are achieved by the understanding and integration of hierarchical visual information and external knowledge proposed in SightSense.
5.5. Ablation Study
To better understand the contribution of each stage in SightSense, we provide a brief ablation study. As demonstrated in Table 3 and Table 2, the performance of the SightSense series improves with the addition of each module, in terms of both explanation similarity and location accuracy. The full version of SightSense outperforms all other simplified versions, validating the important role of hierarchical visual information and external knowledge in the reasoning process. These results indicate that all stages in the proposed SightSense workflow are both effective and necessary.
5.6. Case Study
@@
To further investigate the performance of SightSense and other baseline models, we present a qualitative case study. As shown in Figure 4, SightSense is able to detect the fine-grained visual clues, i.e., Michigan license plate, and perceive the coarse-grained landscape, i.e., abundance of lakes. It then connect such hierarchical visual information with external knowledge to generate the final prediction which is accurate at both country and state level.
In contrast, other baselines demonstrate notable limitation. GPT-4o-mini and DeepSeek generate significant hallucination that are inconsistent with visual evidence. Qwen2VL-7B captures informative clues from the global context but fails to identify detailed visual elements, resulting in a vague location prediction. Lack of integration of hierarchical visual information, these baseline models are unable to generate accurate and logical prediction as SightSense.
6. Conclusion
@@
In this paper, we introduce a novel multimodal reasoning task, GeoGuess, identifing the location of a street view and providing comprehensive explanation. To support GeoGuess, we construct a new benchmark dataset GeoExplain. To our best knowledge, GeoExplain is the first dataset that introduces expert-level explanation on how to infer the location from a given street view image. We further propose a three-stage method SightSense which integrates hierarchical visual information at multiple levels of granularity and external knowledge to solve GeoGuess. Experiments on GeoGuess demonstrates our method achieve the best performance among other state-of-the-art baseline models and provide strong generalizability when transplanting to other models. Moreover, no existing model reach human expert-level accuracy, which validates the difficulty of GeoGuess and its significance as a benchmark toward the development of Artificial General Intelligence.

### v2

@@ 4.3 Visual Clue Detection
$=\cup_{j=1}^{im}\text{OOD}(I_{j},prompt)$ (4)
Based on analysis of high-frequency nouns in the explanation $E$ and knowledge snippets $S$, we design a comprehensive prompt to capture all potentially identifiable object. We use Grounding DINO[37] as our open-vocabulary object detector in this stage.
4.4 Multimodal Knowledge Retrieval
Since GeoGuess is an expert-level task even for humans, relying on the inherent commonsense knowledge in MLLMs obtained from large-scale pretraining, or on reasoning techniques like [71, 77], is insufficient for generating accurate answers. To alleviate potential hallucinations produced by LLM, we have to explicitly incorporate external geographical knowledge into the reasoning process. Accordingly, in the second stage, we aim to retrieve informative knowledge snippets from GeoKnowledge in a multimodal way to assist the reasoning process, inspired by [28].
@@ 4.4 Multimodal Knowledge Retrieval
$\mathbf{S^{*}}$ $=\{S_{i_{1}^{*}},S_{i_{2}^{*}},S_{i_{3}^{*}}\}$ (5)
$\mathbf{V^{*}}$ $=\{V_{j_{1}^{*}},V_{j_{2}^{*}},V_{j_{3}^{*}}\}$ (6)
$i_{1}^{*},i_{2}^{*},i_{3}^{*},j_{1}^{*},j_{2}^{*},j_{3}^{*}$ $=\arg\operatorname{Top\text{-}3}_{i,j}\frac{\text{VE}(KI_{i})\cdot\text{VE}(V_{j})}{|\text{VE}(KI_{i})||\text{VE}(V_{j})|}$ (7)
In Equation 7, $KI$ denotes an image of country-specific object from GeoKnowledge, and $V$ represents detected visual clues obtained in the first stage. To retrieve the top 3 similar image in GeoKnowledge, we employ cosine similarity. A visual encoder, denoted as VE, is used to embed images into vector representations. In our work, we adopt CLIP-ViT [50] as the visual encoder with an embedding dimension of 768.
4.5 Reasoning Generation
@@
After constructing both image and text prompt, we feed them into a pre-trained MLLM to perform reasoning and generate final answer.
4.6 Training Policy
@@
$\text{Loss}=-\sum_{t=1}^{T}P(x_{t}|x_{1},\dots,x_{t-1})$ (8)
5 Experiments
@@
5.1 Experiments Setup
SightSense utilizes GroundingDINO [37] as the open-vocabulary object detector in the visual clue detection stage, CLIP-ViT [50] as the visual encoder in the multimodal knowledge retrieval stage and Qwen2VL-7B-Instruct [68] as the base reasoning model in reasoning generation stage. We finetune only parameters in Qwen2VL and other baselines, while keeping the parameters of models in GroundingDINO and CLIP-ViT frozen. For finetuning details, we adopt Adam optimizer [25] to optimize the training objective and use LinearLR to adjust the learning rate. All training work was completed on NVIDIA H100 PCIe.
5.2 Evaluation Metrics
We refer to the location and explanations generated by SightSense and other baselines as candidate location and explanations. Evaluation of the GeoGuess task comprises two aspects: (1) the accuracy of the candidate location, and (2) the textual similarity between model-generated candidate explanations and human-written references which is measured by both automatic metrics and human evaluation.
5.2.1 Accuracy of Location
We convert the ground-truth geo-coordinate into structured ground-truth location in the format (COUNTRY, STATE, CITY, STREET) which matches the format of candidate location. In case that some countries may have different names, we apply fuzzy match [^2] RapidFuzz and pycountry. Accuracy is computed at each of the four hierarchical levels as Equation 9, where Match is the fuzzy match function, $c_{i}$ and $gt_{i}$ denotes the candidate location and ground-truth location respectively.
$\text{Acc}=\frac{\sum_{i=1}^{n}\text{Match}(c_{i},gt_{i})}{n}$ (9)
5.2.2 Explanation: Automatic Metrics
To measure the similarity between candidate and reference explanations, we employ two aspects of automatic metrics.
  • Relevance: We use (1) BLEU [46], (2) Rogue-L [33], (3) CIDEr [65], (4) Meteor [5] and (5) BERTScore [80] to evaluate the relevance between candidate and reference explanations.
  • Diversity: We use (1) Dist-1 and (2) Dist-2 [29] to measure the diversity of candidate explanations themselves.
@@ 5.2.2 Explanation: Automatic Metrics
P+VC
LLaVA-1.5-7B-VC 3.00 $\uparrow$ 0.90 $\downarrow$ 18.85 $\uparrow$ 1.03 $\downarrow$ 28.29 $\downarrow$ 84.98 $\uparrow$ 5.02 $\downarrow$ 17.70 $\downarrow$
@@
P+VC+K
LLaVA-1.5-7B-KVC 5.44 $\uparrow$ 2.03 $\uparrow$ 20.00 $\uparrow$ 1.12 $\uparrow$ 29.35 $\uparrow$ 85.39 $\uparrow$ 4.58 $\downarrow$ 15.37 $\downarrow$
GPT-4o-mini-KVC 2.45 $\uparrow$ 0.40 $\downarrow$ 13.32 $\uparrow$ 0.37 $\uparrow$ 22.42 $\downarrow$ 82.88 - 7.49 $\uparrow$ 28.16 $\downarrow$
@@
Table 1: Automatic evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines and underlying number denotes the highest accracy for each type. $\uparrow$, $\downarrow$ and - indicates how the accuracy change for corresponding models when introducing more modules.
5.2.3 Explanation: Human Evaluation
@@
5.3 Baselines
We compare SightSense with several state-of-the-art baselines in multimodal reasoning. We validate the effectiveness of hierarchical visual information in the reasoning process even for other baselines by transplanting the first two stages of our pipeline into them. This enables a fair comparison and highlights the generalizability of our approach. Furthermore, we classify the baselines into three different types according to information they receive during generating explanations. Note we do not finetune on models with * due to their close-source nature or inapplicable training policy.
•  Panorama Only: In this type, we use SightSense_wKVC which is a simplified version of SightSense, removing the Visual Clue Detection and Knowledge Retrieval. We selected both geo-localization specified methods and well-established and widely recognized state-of-the-art MLLMs with comparable model sizes: (1) GeoCLIP [66]; (2) StreetCLIP [19]; (3) smileGeo [21]; (4) GeoReasoner [31]; (5) LLaVA-1.5-7B [36]; (6) GPT-4o-mini* and (7) DeepSeek-VL2-tiny [74]. These models serve as strong and representative baselines for evaluating the performance of our approach.
•  Panorama + Visual Clue: Similarly, we finetune SightSense_wK which removes only the Knowledge Retrieval from SightSense. Meanwhile, we transplant the Visual CLue Detection stage into other MLLM baselines. We denote them as (1) LLaVA-1.5-7B-VC; (2)GPT-4o-mini-VC* and (3) DeepSeek-VL2-tiny-VC.
•  Panorama + Visual Clue + Knowledge: We use SightSense with all components enabled. Accordingly, we refer to the baseline models with all three modules as (1) LLaVA-1.5-7B-KVC; (2) GPT-4o-mini-KVC* and (3) DeepSeek-VL2-tiny-KVC.
5.4 Main Results
@@
P+VC
LLaVA-1.5-7B-VC 16.62 $\downarrow$ 1.38 $\uparrow$ 0.00 - 0.00 -
GPT-4o-mini-VC 11.38 $\uparrow$ 0.92 $\uparrow$ 0.00 - 0.00 -
DeepSeek-VL2-tiny-VC 29.79 $\uparrow$ 5.32 $\uparrow$ 0.09 - 0.00 -
@@
P+VC+K
LLaVA-1.5-7B-KVC 25.80 $\uparrow$ 4.91 $\uparrow$ 0.10 $\uparrow$ 0.09 $\uparrow$
@@
Table 2: Accuracy of candidate location in %. Denotations are the same as in Table 1
@@
Table 3: Human Evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines.
@@
Figure 4: A case study of explanation generated by SightSense, LLaVA-1.5-7B, GPT-4o-mini and DeepSeek-VL2-tiny. Given image is a thumbnail of panoramas set and the groundtruth is on the right hand side.
@@
•  Accuracy of Candidate Location: From Table 2, SightSense substantially outperforms all other baselines across all four levels, achieving notable gains of 27% and 18% on the Country and State levels respectively. Meanwhile, within each modality type, the simplified version of SightSense consistently achieves near-best performance except in the Panorama-only type, where SightSense_wKVC underperforms geo-localization-specified models. However, when all components are enabled, SightSense outperforms even these non-explanation geo-localization-specified methods.This verifies the effectiveness of the architecture and training objective of SightSense and highlights its advantage in addressing the GeoGuess task. The slight accuracy drop is due to misleading of confusing visual clues and potentially wrong retrieved knowledge snippets. We observe some illogical inconsistencies happens, such as accuracy drops at state level while they increases at city level. This suggests that finetuning may cause the model to confuse different administrative tiers. In general, the consistent accuracy gains across all four models further verify the robustness and generalizability of our method.
@@
•  Automatic Metric of Candidate Explanation: We observe SightSense series still achieves the best performance regarding explanation similarity evaluation across all types in 1. Explanations generated by SightSense show the highest similarity with the reference explanations. The diversity metric is primarily dominated by DeepSeek-VL2 which means SightSense tends to generate a fixed template learned from the GeoExplain dataset. As with location accuracy analysis, we also notice significant increase in similarity metric when introducing Visual Clue Detection and Multimodal Knowledge Retrieval module, proving the effectiveness of our approach.
@@
•  Human Evaluation of Candidate Explanation: Table 3 demonstrates the result of human evaluation. Except for Fluency, SightSense achieves the highest score across all other metrics. This indicates that the explanation generated by SightSense provide the strongest evidence, the broadest and the most relevant geographical knowledge, and the most complete reasoning chain. These improvements are achieved by the understanding and integration of hierarchical visual information and external knowledge proposed in SightSense.
5.5 Ablation Study
To better understand the contribution of each stage in SightSense, we provide a brief ablation study. From the comparison across the same model in different types, introducing Visual Clue Detection module results in 8 accuracy increases, 2 decreases and 6 ties. Similarly, introducing Multimodal Knowledge Retrieval module leads to 15 increases and 1 decrease. The comparison demonstrates that integrating hierarchical visual information and external knowledge consistently improves location prediction accuracy.
5.6 Case Study
@@
To further investigate the performance of SightSense and other baseline models, we present a qualitative case study. As shown in Figure 4, SightSense is able to detect the fine-grained visual clues, i.e., Michigan license plate, and perceive the coarse-grained landscape, i.e., abundance of lakes. It then connect such hierarchical visual information with external knowledge to generate the final prediction which is accurate at both country and state level. In contrast, other baselines demonstrate notable limitation. GPT-4o-mini and DeepSeek generate significant hallucination that are inconsistent with visual evidence. LLaVA captures informative clues from the global context but fails to identify detailed visual elements, resulting in a vague location prediction.
6 Conclusion
@@
In this paper, we introduce a novel multimodal reasoning task, GeoGuess, identifing the location of a street view and providing comprehensive explanation. To support GeoGuess, we construct a new dataset GeoExplain. To our best knowledge, GeoExplain is the first dataset that introduces expert-level explanation on how to infer the location from a given street view image. We further propose a three-stage method SightSense which understand hierarchical visual information at multiple levels to solve GeoGuess. Experiments on GeoGuess demonstrates our method outperforms other state-of-the-art baseline models and provide strong generalizability when transplanting to other models. Moreover, no existing model reach human-level accuracy, which validates the difficulty of GeoGuess and its significance as a benchmark toward the development of Artificial General Intelligence. Future work could focus on alleviating negative effects of misleading knowledge snippets to further achieve better performance on GeoGuess.